# Constructing a T-test for Value Function Comparison of Individualized Treatment Regimes in the Presence of Multiple Imputation for Missing Data

### v1

@@ 2.4 Comparison Between Value Functions
where the empirical estimates of $T_{0}^{2}$ is:
$\hat{T}_{m}^{2}=m^{-1}\sum_{i=1}^{m}(\hat{U}_{i,1}-\hat{U}_{i,2}-\bar{U}_{m,1}+\bar{U}_{m,2})^{2}+m(\hat{W}_{m,1}-\hat{W}_{m,2})^{T}\hat{\Sigma}_{n}(\hat{W}_{m,1}-\hat{W}_{m,2}).$
@@
$\hat{\Sigma}_{n}=Var(\hat{\theta}_{n})=K^{-1}\sum_{k=1}^{K}VarCov(\hat{\theta}_{n,k})+(1+1/K)(K-1)^{-1}\sum_{k=1}^{K}(\hat{\theta}_{n,k}-\bar{\theta}_{n})^{2},$
where $\bar{\theta}_{n}=1/K\sum_{k=1}^{K}\hat{\theta}_{n,k}$. $\hat{\Sigma}_{n}$ is the same across imputations. It is positive definite because it is the sum of a positive definite matrix and a positive semi-definite matrix.
Thus, the variance for the difference between the value function of two ITRs is:
$Var(\hat{V}_{m}(\hat{d}_{1,n})-\hat{V}_{m}(\hat{d}_{2,n}))=m^{-2}\sum_{i=1}^{m}(\hat{U}_{i,1}-\hat{U}_{i,2}-\bar{U}_{m,1}+\bar{U}_{m,2})^{2}+(\hat{W}_{m,1}-\hat{W}_{m,2})^{T}\hat{\Sigma}_{n}(\hat{W}_{m,1}-\hat{W}_{m,2}).$ (8)
Under the null hypothesis, $V_{0}(\hat{d}_{1,n})-V_{0}(\hat{d}_{2,n})=0$. The t-test statistics $t$ for this null hypothesis can be constructed as follows:
$t=\frac{\hat{V}_{m}(\hat{d}_{1,n})-\hat{V}_{m}(\hat{d}_{2,n})}{T_{m}/\sqrt{m}}\sim N(0,1)$ (9)
The p-value for this t-test can be obtained by the test statistic $t$.
@@ 2.5 Multiple Imputation
In the presence of missing data, we extend our method to address the variance of the value function in the case of multiple imputation. This extension is based on the ideas from Chapter 2.3.2 from Van Buuren (2018) Van Buuren (2018). Suppose we have K imputations. Let $k=1,\dots,K$ be the index of the multiple-imputed data. Let $\hat{V}_{m,k}(\hat{d}_{j})$ denote the value function for the $k^{th}$ imputed test set. Let $\hat{\sigma}_{m,j,k}=Var(\hat{V}_{m,k}(\hat{d}_{j}))$ be the estimator of the variance-covariance matrix for the estimated value function for the $k^{th}$ imputed test set. Let $\hat{T}^{2}_{m,k}=Var(\hat{V}_{m,k}(\hat{d}_{1})-\hat{V}_{m,k}(\hat{d}_{2}))$ denote the variance for the difference between two value functions for the $k^{th}$ imputed test set. The subscript m indicates that the estimates are obtained for the test set.
Let $\tilde{V}_{m}(\hat{d}_{j})=K^{-1}\sum_{k=1}^{K}\hat{V}_{m,k}(\hat{d}_{j})$, be value function estimate for decision rule $\hat{d}_{j}$. Let $\sigma_{0,K}^{2}=Var_{K}(\hat{V}(\hat{d}_{j}))$ denote the variance associated with the estimate for a single value function. The empirical estimate of $\sigma_{0,K}^{2}$ is:
@@
For the paired test, Let $T_{0,K}^{2}=Var_{K}(\hat{V}(\hat{d}_{1})-\hat{V}(\hat{d}_{2}))$ denote the variance for the difference between value functions for decision rule $\hat{d}_{1}$ and $\hat{d}_{2}$. The empirical estimate of $T_{0,K}^{2}$ is:
@@
Under the null hypothesis, $V_{0,K}(\hat{d}_{1,n})-V_{0,K}(\hat{d}_{2,n})=0$. The t-test statistics $t_{K}$ for this null hypothesis can be constructed as follows:
$t_{K}=\frac{K^{-1}\sum_{k=1}^{K}(\hat{V}_{m,k}(\hat{d}_{1,n})-\hat{V}_{k,m}(\hat{d}_{2,n}))}{\sqrt{Var_{K}(\hat{V}(\hat{d}_{1})-\hat{V}(\hat{d}_{2}))}}\sim N(0,1)$ (12)
@@ 3 Simulation
Figure 1: The value function results of four simulation scenarios: (a) with random treatment assignment and complete data, (b) with random treatment assignment and data missing at random (c) with treatment assignment depending on the covariates and complete data, (d) with treatment assignment depending on the covariates and data missing at random. Within each scenario, we compare the value functions of five different treatment regimes which are indicated in the x-axis: observed treatment, treatment $A=0$ for all individuals, treatment $A=1$ for all individuals, Q-learning optimal ITR, and D-learning optimal ITR. For comparison purposes, we include the average outcome as a benchmark. The true value function is presented with the grey line and the variation of the true value function across the 10 replicates is indicated by the grey shaded area.
@@
Figure 2: The differences of value functions between selected pairs of the following treatments: observed treatment, universal treatment $A=0$, universal treatment $A=1$, Q-learning’s optimal ITR, and D-learning’s optimal ITR. The results are presented under four simulation scenarios.
@@
We evaluate and compare the value functions obtained from applying each of the following treatment regimes to the population under two scenarios: (1) the observed treatment, (2-3) the one-size-fits-all approach (one for each of the binary treatments), (4) Q-learning ITR, (5) D-learning ITR. A convex penalty, elastic net penalization (Zou and Hastie, 2005), was used for each of the ITR models for variable selection. We generate a total of 5000 observations for each simulated sample with 70% for the training set and 30% for the test set. The number of covariates included in the model is $p=20$. We perform 10 replicates for each of the simulations.
@@
The following data generation process is modified from the first simulation scenario in Tian et al. (2014). The covariates $X$ are uncorrelated and generated from a multivariate normal distribution with mean zero and the covariance-variance matrix is a diagonal of 1s. Let $A$ denote the binary treatment. In the model implementation step, we will use $A=\{0,1\}$ for Q-learning and switch to $A=\{-1,1\}$ for D-learning. Let $Y=\{0,1\}$ denote the binary outcome. $\varepsilon\sim N(0,1)$ is the random error for generating the outcome $Y$. Let $\beta_{0}=(\sqrt{6})^{-1}$, $\beta_{1}=\beta_{2}=0$, $\beta_{j}=(2\sqrt{6})^{-1}$, $j=3,4,\dots,10$, $\beta_{11}=...=\beta_{p}=0$ be the coefficient for the main effect. Let $(\gamma_{0},\gamma_{1},\gamma_{2},\gamma_{3},\gamma_{4},\gamma_{5},\dots,\gamma_{p})=(1,1,-1,1,-1,0,\dots,0)$ be the coefficients for treatment interactions. Let $Y\sim binomial$ with probability equal to the expit of
$(\beta_{0}+\sum_{j=1}^{p}\beta_{j}X_{j})^{2}+(\gamma_{0}+\sum_{j=1}^{p}\gamma_{j}X_{j})A+\varepsilon.$
@@
The generated data is designed to have small main effects $(\beta_{0},\dots,\beta_{p})$, a large treatment effect $\gamma_{0}$, and large treatment interaction effects $(\gamma_{1},\dots,\gamma_{p})$. To align with the application example presented in section 4, we assume, without loss of generality, that a smaller outcome Y and a smaller value function are preferable. We consider four scenarios:
@@
In scenarios where data are missing at random, the missingness mechanism is implemented by assigning a conditional probability of missingness to the variables $X_{2},\ldots,X_{p}$. This probability is contingent upon the value of $X_{1}$: specifically, there is a 15% probability of missingness when $X_{1}>0$, and a notably lower probability of 10% in cases where this condition is not met. Under each scenario, the true value function is estimated based on a simulated large test set with 10000 individuals. The data generation process for this test set is the same as for the training set under the same scenario. The true variance for the value function is $\hat{\tau}_{j}^{2}=\frac{1}{L}\sum_{l=1}^{L}(V_{l}(\hat{d}_{j})-V_{*}(\hat{d}_{j}))^{2}$. $V_{l}(\hat{d}_{j})$ is the estimated value function based on the test set $l$ where we plug in the estimated decision rule $\hat{d}_{j}$ and the estimated propensity score model from the training set. $V_{*}(\hat{d}_{j})$ is the estimated value function based on the giant test set, and is based on the true propensity score parameters and the decision rule $\hat{d}_{j}$ estimated from the training set. The true variance for the difference between value functions is constructed similarly.
@@
Figure 1 shows the value functions across these four scenarios. We observe that the value function for ITRs from Q-learning and D-learning are the lowest among all methods. The value function estimates associated with the optimal ITRs estimated using Q-learning and D-learning are close to the true value function. The estimated variance associated with these two value functions is also close to the true variance around the true value function. Since Q-learning shows the best value function and lowest optimal treatment misclassification rate overall (Table 1), we focus on comparing Q-learning with the others in the pairwise comparison as shown in Figure 2. We observe that Q-learning is significantly better than the one-size-fits-all treatments ($A=0$ and $A=1$) in all four scenarios. This is expected because we simulate the data with a large prescriptive treatment effect. We observe that the difference is large between value functions for assigning treatment $A=0$ and $A=1$ to all individuals. This is expected because we simulate the data with a large treatment effect. For observational data, when treatment assignments are nonrandom and probabilities need to be estimated by propensity scores, value function estimates have larger variances compared to the value function estimates for randomized trial data (Table 1). For datasets with missing values, the performance remains comparable to that of complete datasets when the missing data is random and is addressed through multiple imputation using missForest. The results shown here are the test set results. The training set results are similar to the test set results but with a smaller variance. The simulation results illustrate the effectiveness of our approach in discerning disparities between two ITRs while maintaining reasonable variance estimates in both observational data and randomized trial data, with or without missing values. Additionally, it highlights a constraint of our method that the accuracy of the estimate’s variance relies on the precision of propensity score estimation.
@@
Table 1: Estimated value functions and treatment misclassification rate (MC) of different treatment regimes and under four scenarios for simulation data.
@@
Mean(SD) MC Mean(SD) MC Mean(SD) MC Mean(SD) MC
Obs 0.631(0.026) 0.509 0.631(0.026) 0.509 0.631(0.028) 0.581 0.631(0.027) 0.581
All0 0.567(0.018) 0.235 0.567(0.018) 0.235 0.561(0.020) 0.235 0.562(0.019) 0.235
All1 0.696(0.017) 0.765 0.696(0.017) 0.765 0.699(0.020) 0.765 0.700(0.019) 0.765
Q-Learning 0.534(0.019) 0.064 0.535(0.019) 0.065 0.529(0.022) 0.063 0.532(0.021) 0.066
D-Learning 0.536(0.019) 0.087 0.537(0.019) 0.089 0.534(0.022) 0.094 0.535(0.021) 0.095

### v3

@@ 2.4 Comparison Between Value Functions
where
$T_{0}^{2}=Var(U_{1}-U_{2})+\frac{m}{n}(W_{1}-W_{2})^{T}\Sigma_{0}(W_{1}-W_{2}).$
Thus, the variance for the difference between the value function of two ITRs is approximately:
$Var(\hat{V}_{m}(\hat{d}_{1,n})-\hat{V}_{m}(\hat{d}_{2,n}))=m^{-2}\sum_{i=1}^{m}(\hat{U}_{i,1}-\hat{U}_{i,2}-\bar{U}_{1}+\bar{U}_{2})^{2}+\frac{1}{n}(\hat{W}_{1}-\hat{W}_{2})^{T}\hat{\Sigma}_{n}(\hat{W}_{1}-\hat{W}_{2}).$ (11)
@@
$\hat{\Sigma}_{n}=nVar(\hat{\theta}_{n})=nK^{-1}\sum_{k=1}^{K}Var(\hat{\theta}_{n,k})+n(1+1/K)(K-1)^{-1}\sum_{k=1}^{K}(\hat{\theta}_{n,k}-\bar{\theta}_{n})^{2},$
where $\bar{\theta}_{n}=1/K\sum_{k=1}^{K}\hat{\theta}_{n,k}$. $\hat{\Sigma}_{n}$ is the same across imputations, and it is positive definite because it is the sum of a positive definite matrix and a positive semi-definite matrix. Under the null hypothesis, $V_{0}(\hat{d}_{1,n})-V_{0}(\hat{d}_{2,n})=0$. The t-test statistics $t$ for this null hypothesis can then be constructed as follows:
$t=\frac{\hat{V}_{m}(\hat{d}_{1,n})-\hat{V}_{m}(\hat{d}_{2,n})}{\sqrt{Var(\hat{V}_{m}(\hat{d}_{1,n})-\hat{V}_{m}(\hat{d}_{2,n}))}}\sim N(0,1),$ (12)
and this can be used to compute the p-value.
@@ 2.5 Multiple Imputation
In the presence of missing data, we extend our method to address the variance of the value function in the case of multiple imputation. This extension is based on the ideas from Chapter 2.3.2 from Van Buuren (2018) 17. Suppose we have K imputations. Let $k=1,\dots,K$ be the index of the multiple-imputed data. The subscript m indicates that the estimates are obtained for the test set.
For the single decision rule $\hat{d}_{j}$, let $\hat{V}_{m,k}(\hat{d}_{j})$ denote the estimated value function for the $k^{th}$ imputed test set. Let $\tilde{V}_{m}(\hat{d}_{j})=K^{-1}\sum_{k=1}^{K}\hat{V}_{m,k}(\hat{d}_{j})$, be value function estimate for decision rule $\hat{d}_{j}$. Let $\hat{\sigma}_{m,j,k}=Var(\hat{V}_{m,k}(\hat{d}_{j}))$ be the estimator of the variance-covariance matrix for the estimated value function for the $k^{th}$ imputed test set. Let $\sigma_{0,K}^{2}=Var_{K}(\hat{V}(\hat{d}_{j}))$ denote the variance associated with the estimate for a single value function. The empirical estimate of $\sigma_{0,K}^{2}$ is:
@@
For the paired test between decision rules $\hat{d}_{1}$ and $\hat{d}_{2}$, let $\hat{T}^{2}_{m,k}=Var(\hat{V}_{m,k}(\hat{d}_{1})-\hat{V}_{m,k}(\hat{d}_{2}))$ denote the estimated variance for the difference between two value functions for the $k^{th}$ imputed test set. Let $T_{0,K}^{2}=Var_{K}(\hat{V}(\hat{d}_{1})-\hat{V}(\hat{d}_{2}))$ denote the variance for the difference between value functions for decision rule $\hat{d}_{1}$ and $\hat{d}_{2}$. The empirical estimate of $T_{0,K}^{2}$ is:
@@
Under the null hypothesis, $H_{0}:V_{0,K}(\hat{d}_{1,n})-V_{0,K}(\hat{d}_{2,n})=0$. The t-test statistics $t_{K}$ for this null hypothesis can be constructed as follows:
$t_{K}=\frac{K^{-1}\sum_{k=1}^{K}(\hat{V}_{m,k}(\hat{d}_{1,n})-\hat{V}_{m,k}(\hat{d}_{2,n}))}{\sqrt{Var_{K}(\hat{V}_{m,K}(\hat{d}_{1,n})-\hat{V}_{m,K}(\hat{d}_{2,n}))}}\sim N(0,1)$ (15)
@@ 3 Simulation
Figure 1: The value function results of four simulation scenarios with $\delta_{t}=1$ and $n=5000$: (a) with random treatment assignment and complete data, (b) with random treatment assignment and data missing at random (c) with treatment assignment depending on the covariates and complete data, (d) with treatment assignment depending on the covariates and data missing at random. Within each scenario, we compare the value functions of eight different treatment regimes which are indicated in the x-axis: observed treatment, treatment $A=0$ for all individuals, treatment $A=1$ for all individuals, Q-learning optimal ITR, and D-learning optimal ITR, treatment rule $d_{1}$, treatment rule $d_{2}$, and the true optimal ITR. The true value function is presented with the grey line. The estimated values and 95% confidence intervals from our method, bootstrap estimates, and the true empirical 95% confidence intervals are depicted using varying shades of gray.
@@
Figure 2: The differences of value functions and the associate standard deviation between selected pairs of the following treatments: observed treatment, universal treatment $A=0$, universal treatment $A=1$, Q-learning’s optimal ITR, D-learning’s optimal ITR, treatment rule $d_{1}$, treatment rule $d_{2}$, and the true optimal ITR. The results are presented under four simulation scenarios with $\delta_{t}=1$ and $n=5000$, comparing our estimates, bootstrap estimates, and the true values and 95% confidence intervals.
@@
We evaluate and compare the value functions obtained from applying each of the following treatment regimes to the population under two scenarios: (1) the observed treatment, (2-3) the one-size-fits-all approach (one for each of the binary treatments), (4) Q-learning ITR, (5) D-learning ITR. A convex penalty, elastic net penalization 18, was used for each of the ITR models for variable selection. We generate a total of 5000 observations for each simulated training set and 1500 observations for each test set. The number of covariates included in the model is $p=20$. We perform 1000 replicates for each of the simulations.
@@
The following data generation process is modified from the first simulation scenario in Tian et al 5. The covariates $X$ are generated from three different underlying distributions: $X_{1},...,X_{10}$ are uncorrelated and generated from a multivariate normal distribution with mean zero and the covariance-variance matrix is a diagonal of 1s; $X_{11},...,X_{15}$ are generated from a uniform distribution $U(-0.5,0.5)$; $X_{16},...,X_{20}$ are generated from a binary distribution with $p=0.5$. Let $A$ denote the binary treatment. In the model implementation step, we will use $A=\{0,1\}$ for Q-learning and switch to $A=\{-1,1\}$ for D-learning. $\varepsilon\sim N(0,1)$ is the random error [^2] The error term has been incorporated into the outcome data generation mechanism to better reflect potential real-world inaccuracies. for generating the outcome $Y$. Let $\beta_{0}=(\sqrt{6})^{-1}$, $\beta_{1}=\beta_{2}=0$, $\beta_{j}=(2\sqrt{6})^{-1}$, $j=3,4,\dots,10$, $\beta_{11}=...=\beta_{p}=0$ be the coefficient for the main effect. Let $Y\sim binomial$ with probability equal to the expit of
$(\beta_{0}+\sum_{j=1}^{p}\beta_{j}X_{j})^{2}+(I(X_{1}>0)+I(X_{2}>0)-2I(X_{1}\leq 0)I(X_{2}\leq 0))A\delta_{t}+\varepsilon.$
@@
The generated data is designed to have main effects $(\beta_{0},\dots,\beta_{p})$, and treatment interaction effect $(I(X_{1}>0)+I(X_{2}>0)-2I(X_{1}\leq 0)I(X_{2}\leq 0))\delta_{t}$, where $\delta_{t}>0$ is a parameter that controls the magnitude of the treatment interaction effect on the outcome. To align with the application example presented in section 4, we assume, without loss of generality, that a smaller outcome Y and a smaller value function are preferable. Thus, the optimal treatment regime is $d_{opt}=I(I(X_{1}>0)+I(X_{2}>0)-2I(X_{1}\leq 0)I(X_{2}\leq 0))\leq 0=I(X_{1}\leq 0\,and\,X_{2}\leq 0)$. We consider four scenarios:
@@
In scenarios where data are missing at random, the missingness mechanism is implemented by assigning a conditional probability of missingness to the variables $X_{2},\ldots,X_{p}$. This probability is contingent upon the value of $X_{1}$: specifically, there is a 15% probability of missingness when $X_{1}>0$, and a lower probability of 10% in cases where this condition is not met.
The true value function $V_{*}(\hat{d}_{j})$ is obtained from applying each decision rule $\hat{d}_{j}$ on a simulated large test set comprising 100,000 individuals. The data generation process for this large test set mirrors that of the corresponding complete training set. The true value function $V_{*}(\hat{d}_{j})$ is calculated based on the true propensity score parameters and the decision rule $\hat{d}_{j}$ derived from the training set. We define the true variance of the value function as $\hat{\tau}_{j}^{2}=\frac{1}{L}\sum_{l=1}^{L}(V_{l}(\hat{d}_{j})-V_{*}(\hat{d}_{j}))^{2}$, where $V_{l}(\hat{d}_{j})$ represents the estimated value function from test set $l$, utilizing the estimated decision rule $\hat{d}_{j}$ and the propensity score model from the corresponding training set. The true variance for the difference between value functions is constructed in a similar manner.
The point estimates for value functions are derived by averaging the value function estimates across 1000 replications on the test set. The estimated variance for each value function is computed using Equation 9, while the variances for differences between two value functions are calculated using Equations 11 (for scenarios 1 and 3) and 14 (for scenarios 2 and 4), also averaged over 1000 replications. For bootstrapping estimates, the point estimate of the value function is the mean obtained from 1000 bootstrap samples (we have also obtained estimates from 100, 200, 500 bootstrapping samples, the results are similar to 1000 bootstraping samples and are thus omited here), which is then further averaged across 1000 replications. Similarly, the bootstrapping variance is determined by calculating the variance of the 1000 bootstrap samples and averaging this across 1000 replications.
@@
Figure 1 shows the value functions and standard deviation in four scenarios (a)-(d). We observe that the point estimates of value functions using our estimated method are close to that estimated from bootstrapping and the true value, across all 4 scenarios and almost all methods. The variance estimates of a single value function using our estimated method are close to that estimated from bootstrapping and the true value (except that our method gives a slightly larger variance estimate for the observed treatment). Figure 2 shows that, for comparing pairwise value functions, the point estimate and the variance estimated for the difference from our method are also close to the bootstrapping estimates and the truth. As expected, the value function for ITRs from $d_{opt}$ is the lowest among all methods. Table 1 presents the difference, standard deviation (SD), and t-test rejection rates for comparisons between $d_{1}$ vs. $d_{2}$ and $d_{1}$ vs. $d_{opt}$, under the setting $\delta_{t}=1$ and $n=5000$. The point estimates and SDs are consistent across our estimates, bootstrapped estimates, and the true value differences. The t-test for the difference between $d_{1}$ and $d_{2}$ yields a rejection rate close to zero, whereas the test for the difference between $d_{1}$ and $d_{opt}$ results in a rejection rate close to one. For observational data, when treatment assignments are nonrandom and probabilities need to be estimated by propensity scores, value function estimates have slightly larger variances compared to the value function estimates for randomized trial data (Table 1). For datasets with missing values, the performance remains comparable to that of complete datasets when the missing data is random and is addressed through multiple imputation using missForest. The results shown here are the test set results. The training set results are similar to the test set results but with a smaller variance. The simulation results illustrate the effectiveness of our approach in discerning disparities between two ITRs while maintaining reasonable variance estimates in both observational data and randomized trial data, with or without missing values.
@@
Table 1: Estimated difference in value functions Mean (SD) and t-test rejection rate (RR) under four scenarios for simulation data based on 1000 replications ($\delta_{t}=1$, $n=5000$).
@@
Method Test Mean(SD) RR Mean(SD) RR Mean(SD) RR Mean(SD) RR
Est $d_{1}$ vs. $d_{2}$0.001(0.018) 0.069 -0.004(0.018) 0.089 0.000(0.021) 0.039 -0.004(0.021) 0.055
Boot $d_{1}$ vs. $d_{2}$0.001(0.018) 0.069 -0.004(0.018) 0.085 0.000(0.021) 0.041 -0.004(0.020) 0.062
True $d_{1}$ vs. $d_{2}$ 0.001(0.019) 0.000 0.001(0.021) $0$0.002(0.020) $0$0.002(0.022) 0.000
Est $d_{1}$ vs. $d_{opt}$ 0.046(0.013) 0.940 0.043(0.013) 0.896 0.046(0.015) 0.861 0.044(0.015) 0.808
Boot $d_{1}$ vs. $d_{opt}$ 0.046(0.013) 0.939 0.043(0.013) 0.900 0.045(0.015) 0.869 0.044(0.015) 0.822
True $d_{1}$ vs. $d_{opt}$ 0.047(0.013) 1.000 0.047(0.014) 1.000 0.047(0.015) 1.000 0.047(0.016) 1.000
For our power analysis, we define two regimes: $d_{1}=I(X_{1}\leq 0)$, $d_{2}=I(X_{2}\leq 0)$. By design, $V_{0}(\hat{d}_{1,n})=V_{0}(\hat{d}_{2,n})$ and $V_{0}(\hat{d}_{1,n})\neq V_{0}(\hat{d}_{opt,n})$. The power analysis includes different training sample sizes $n\in\{1000,2000,3500,5000\}$ and varying values of $\delta_{t}\in\{1,2,5,10\}$, where treatment interaction hyperparameter $\delta_{t}$ controls the magnitude of treatment heterogeneity. As $\delta_{t}$ increases, the difference between value functions of $d_{1}$ and $d_{opt}$ also increases. Table 2 show the Type I Error rate by testing $H_{0}:d_{1}=d_{2}$. Table 3 shows the power (or 1-Type II Error rate) by testing $H_{0}:d_{1}=d_{opt}$. These results show that both the Type I Errors and the power of our method are slightly lower or similar to that of bootstrapping method across 4 scenarios, across different effect size and different sample sizes. [^3] The results for $\delta_{t}\in\{5,10\}$ were similar, we only present the results for $\delta_{t}\in\{1,2\}$ for clarity.
Table 2: Type 1 Error by sample size and $\delta_{t}$.
Scenario 1 Scenario 2 Scenario 3 Scenario 4
n Method $\delta_{t}$ = 1 $\delta_{t}$ = 2 $\delta_{t}$ = 1 $\delta_{t}$ = 2 $\delta_{t}$ = 1 $\delta_{t}$ = 2 $\delta_{t}$ = 1 $\delta_{t}$ = 2
1000 Est 0.041 0.054 0.061 0.070 0.038 0.037 0.052 0.041
1000 Boot 0.047 0.060 0.066 0.075 0.060 0.075 0.077 0.086
2000 Est 0.053 0.057 0.062 0.087 0.049 0.044 0.062 0.063
2000 Boot 0.058 0.063 0.069 0.090 0.059 0.066 0.079 0.084
3500 Est 0.059 0.064 0.086 0.085 0.049 0.051 0.057 0.067
3500 Boot 0.062 0.064 0.085 0.085 0.056 0.057 0.067 0.076
5000 Est 0.069 0.063 0.089 0.097 0.039 0.039 0.055 0.071
5000 Boot 0.069 0.065 0.085 0.097 0.041 0.047 0.062 0.076
Table 3: Power by sample size and $\delta_{t}$.
Scenario 1 Scenario 2 Scenario 3 Scenario 4
n Method $\delta_{t}$ = 1 $\delta_{t}$ = 2 $\delta_{t}$ = 1 $\delta_{t}$ = 2 $\delta_{t}$ = 1 $\delta_{t}$ = 2 $\delta_{t}$ = 1 $\delta_{t}$ = 2
1000 Est 0.916 1.000 0.850 0.997 0.811 0.998 0.757 0.996
1000 Boot 0.925 1.000 0.860 0.998 0.879 0.999 0.838 0.999
2000 Est 0.925 0.999 0.870 0.999 0.876 1.000 0.833 1.000
2000 Boot 0.929 0.999 0.875 0.999 0.899 1.000 0.869 1.000
3500 Est 0.918 1.000 0.864 1.000 0.880 0.999 0.846 0.998
3500 Boot 0.917 1.000 0.866 1.000 0.892 0.999 0.866 0.999
5000 Est 0.940 1.000 0.896 1.000 0.861 1.000 0.808 1.000
5000 Boot 0.939 1.000 0.900 1.000 0.869 1.000 0.822 1.000
Since our method relies on the correct specification of the propensity model, we examine how propensity score estimation influences the value function and standard deviation estimates (Figure S1) as well as the differences between value functions and their associated standard deviation estimates (Figure S2). The sensitivity analysis indicates that the results remain similar whether using the estimated propensity score or the true propensity score. However, our method may exhibit bias when the propensity score is misspecified; however, the extent of this bias is comparable to that observed in the bootstrap estimates. These findings highlight a limitation of our approach: the accuracy of variance estimation depends on the precision of propensity score estimation. Additionally, they provide insights into how our method performs under model misspecification in certain conditions (Figure S3 illustrates the distribution of selected propensity score models in relation to the true propensity score across the these four models).